A sample of gas has a mass of 0.565  g . Its volume is 115  mL at a temperature of 85  ∘C and a pressure of 760  Torr .

Part A

Find the molar mass of the gas.
Answer #1

PV = nRT

where, P = pressure = 760 torr = 1 atm

V = volume = 0.115 L

n = number of moles

R = Gas constant

T = temperature = 85 + 273 = 358 K

1 * 0.115 = n * 0.0821 * 358

0.115 = n * 29.4

n = 0.115 / 29.4 = 0.00391 mole

number of moles = mass of gas / molar mass of gas

0.00391 mole = 0.565 g / molar mass of gas

molar mass of gas = 0.565 g / 0.0039 mole = 145 g/mol

Therefore, molar mass of gas = 145 g/mol